iam not understanding the question and in fact i have no idea how to solve it! plz help
let f(x)={ (x+2p if x<-2) }
{ (3px+q if x -2<=x<=1) }
{ (3x-2q if x>1) }
determine the values of p and q that make f continuous

Answer by ewatrrr(24785)

to be continuous..need to look at when x = -2 and x-1
when x = -2
x + 2p = 3px + q
Substitute x = -2
-2 + 2p = -6p + q
-2 = -8p + q ****
when x = 1
3px + q = 3x - 2q
Substitute x = 1
3p + q = 3 - 2q
3p + 3q = 3 ***
____________________
-8p + q = -2
3p +3q = 3
__________________ Multiplying 1st by -3 to eliminate q when adding both
27p = 9
p = 1/3
q = 2/3